OpenStudy (anonymous):

The table below shows the values of f(n) for different values of n: n 1 2 3 4 5 6 f(n) 1 2 5 12 29 70 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = 2f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = 2f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

@ganeshie8 plz help

OpenStudy (muzzack):

do you know how to do a sequence? @hero a lil help

OpenStudy (muzzack):

well i know that we replace the number oh wait... Arithmetic Recursive Formula Examine the recursive formula given the arithmetic sequence 2, 5, 8, 11, … The first term of the sequence is 2, and the common difference is 3. f(1) = 2, d = 3 To find the next term in the sequence, the common difference is added to the previous term. f(2) = f(1) + d = 2 + 3 = 5 f(3) = f(2) + d = 5 + 3 = 8 To find the nth term of this sequence, the common difference would be added to the term before it. This is the generic form for an arithmetic recursive formula. f(n) = f(n – 1) + d The variable d is the common difference and n > 0. Check to see that this works by finding the fourth term of the sequence. (n = 4)

OpenStudy (anonymous):

is it a @muzzack ?

OpenStudy (anonymous):

@ganeshie8 plz help

OpenStudy (muzzack):

well what do you think??

OpenStudy (anonymous):

no idea lol

OpenStudy (muzzack):

well believe ur gut

OpenStudy (amistre64):

which rule fits?

OpenStudy (anonymous):

idk lol

OpenStudy (amistre64):

so the question becomes .. why dont you know? you have 4 options to test out, and they give you the values ....

OpenStudy (anonymous):

what are the values? what are you supposed to do with the options? (you can see im not really good at this stuff)

OpenStudy (amistre64):

the values are listed for you already ... for exmaple, when n=3, f(3)=5 you have to determine which rule best defines the function.

OpenStudy (amistre64):

notice that there is no n=0 .. so f(n-3) means nothing to us when n=3. what does this tell us about options with f(n-3) in them?

OpenStudy (anonymous):

tht f(0)=0 i think?

OpenStudy (amistre64):

f(0) is not defined.

OpenStudy (anonymous):

i meant tht lol

OpenStudy (anonymous):

so i have to do f(?)

OpenStudy (amistre64):

so, we can rule out any option that has f(n-3) in it since they all say n starts at 3, and f(3-3) is simply undefined.

OpenStudy (anonymous):

So there is only a and c left over

OpenStudy (amistre64):

right, so now lets follow the rules defined: f(n) = 2f(n -1) f(n - 2); n > 2 , so let n=3 f(3) = 2f(3 -1) f(3 - 2) f(3) = 2f(2) f(1) , is this true? --------------------------------- f(n) = 2f(n -1) + f(n - 2) f(3) = 2f(3 -1) + f(3 - 2) f(3) = 2f(2) + f(1), or is this true?

OpenStudy (anonymous):

hold on i have to substitute the values in the n right? give me a few mins to do tht

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

The second one right @amistre64 ?

OpenStudy (amistre64):

yes: 5 = 2(2)+1 is true.

OpenStudy (anonymous):

Thank You!

OpenStudy (amistre64):

now lets see if it works for the next value, im sure it does, 12 = 5(2) + 2 ... seems to work fine